# Re-parameterized Low-rank Prompt: Generalize a Vision-Language Model within 0.5K Parameters

### v2

@@ Abstract
With the development of large pre-trained vision-language models, how to effectively transfer the knowledge of such foundational models to downstream tasks becomes a hot topic, especially in a data-deficient scenario. Recently, prompt tuning has become a popular solution. When adapting the vision-language models, researchers freeze the parameters in the backbone and only design and tune the prompts. On the one hand, the delicate design of prompt tuning exhibits strong performance. On the other hand, complicated structures and update rules largely increase the computation and storage cost. Motivated by the observation that the evolution pattern of the generalization capability in visual-language models aligns harmoniously with the trend of rank variations in the prompt matrix during adaptation, we design a new type of prompt, Re-parameterized Low-rank Prompt (RLP), for both efficient and effective adaptation. Our method could largely reduce the number of tunable parameters and storage space, which is quite beneficial in resource-limited scenarios. Extensive experiments further demonstrate the superiority of RLP. In particular, RLP shows comparable or even stronger performance than the latest state-of-the-art methods with an extremely small number of parameters. On a series of tasks over 11 datasets, RLP significantly increases the average downstream accuracy of classic prompt tuning by up to 5.25% using merely 0.5K parameters.
@@ 1 Introduction
Figure 1: Comparison between classic prompt tuning and Re-parameterized Low-rank Prompt (RLP) tuning. RLP introduces two little parameter matrices in shape $[n,r]$ and $[r,d]$ separately, and uses the product of them as an equivalent low-rank prompt in shape $[n,d]$.
However, traditional fine-tuning is not a good choice to adapt vision-language models. Simply fine-tuning all the parameters can easily cause the model to forget task-agnostic knowledge that is general and important, and the huge training and storage cost is also an intractable problem. Prompt engineering, a recent popular method in natural language processing and computer vision, provides an overwhelming solution for the adaptation of vision-language models. Prompts are virtual tokens that are manually appended or prepended to the semantic tokens. Such virtual tokens can be hand-crafted designed and frozen [41], and they can be soft and tunable as well [30, 27, 18] for better transferability on different tasks. Typical prompt tuning methods freeze all the other parameters in the model and merely tune the prompts during adaptation.
In the vision-language field, prompt-based methods show good performance as well. CLIP [37] uses a hand-crafted text template for better zero-shot learning. Bahng et al. [1] adopts prompt tuning merely on the image encoder, CoOp [54] uses tunable text prompts to replace the fixed template in CLIP, CoCoOp [55] utilizes image feature to instruct the optimization of the soft text prompt in CoOp.  [21, 24] simultaneously optimize image and text prompts and establish extra connections between isolated modalities.  [22, 47, 3, 53] integrate strong regularization modules or losses into prompt tuning to diminish the overfitting and catastrophic forgetting problem. On the one hand, such delicate designs of prompt tuning in recent methods exhibits strong performance. On the other hand, their complicated structures and update rules largely increase the computation and storage cost.
Motivated by the observation that the evolution pattern of the generalization capability in visual-language models aligns harmoniously with the trend of rank variations in the prompt matrix during adaptation as shown in Figs. 2 and 1, we propose Re-parameterized Low-rank Prompt (RLP) to reach effective and efficient adaptation for vision-language models. For more details about the motivation and our observation, please refer to Sec. 2 and we will provide a thorough analysis. In this paper, we have explored how to effectively adapt vision-language models under an extremely small number of parameters. By choosing a small rank $r$, we could exceedingly compress the parameters used in adaptation. To the best of our knowledge, we are the first the reach comparable or even better performance compared with the latest state-of-the-art tuning methods in the base-to-new generalization and domain generalization using such barely few parameters, i.e. 0.5K.
Additionally, though more complex and delicate, existing methods are still refined on a common fundamental basis, i.e. prompt tuning. Such a common basis guarantees that our RLP could be easily and smoothly integrated into most of the off-the-shelf methods besides individually applied.
@@
  • Motivated by the observation that the evolution pattern of the generalization capability in visual-language models aligns harmoniously with the trend of rank variations in the prompt matrix during adaptation, we propose to re-parameterize classic prompts into low-rank prompts for better accuracy and efficiency.
  • We propose a novel initialization method and integrate Dropout layer as a lightweight regularization module to further improve the performance of low-rank prompt tuning without introducing any extra parameters and inference cost.
@@
  • We conduct extensive experiments and show the fantastic effectiveness and efficiency of our proposed method. In base-to-new generalization, domain generalization, cross-dataset transfer and few-shot learning settings, RLP consistently gives very competitive results, surpassing many state-of-the-art tuning methods though we use fewer parameters.
2 Motivation for Low-Rank Prompts
In this section, we provide the motivation for our design of low-rank prompts. First of all, we do classic prompt tuning on a CLIP [37] model following CoOp [54] and save checkpoint for every epoch. The prompts are randomly initialized here for clear observation. Since we want to inspect the rank properties of the prompt matrix, we do singular value decomposition (SVD) for the prompt matrix $P\in\mathbb{R}^{n\times d}$.
$P=U\Sigma V^{T}$ (1)
where $\Sigma$ is the singular matrix and the values on its main diagonal denoted by $\sigma_{i}$ $(1\leq i\leq m)$ are singular values, which are non-negative. Here $m=min(n,d)$. The number of non-zero singular values is exactly the rank of the prompt matrix. In practice, even though all of the singular values are non-zero, the original matrix could be approximately considered a low-rank matrix as well if some singular values are far smaller than others. Therefore, we compute two values, ``std'' and ``scale'', to investigate the low-rank property of singular values in training time. At a certain point during training, we denote the largest singular value as $\sigma_{max}$ while the smallest singular value as $\sigma_{min}$. Since the SVD result is not unique, we do a scaling transformation first and then compute the standard deviation of the singular values.
$\forall 1\leq i\leq m,\sigma_{i}^{\prime}=\frac{\sigma_{i}}{\sigma_{min}}$ (2)
$\mbox{std}=\sqrt{\frac{1}{m}\sum_{i=1}^{m}{(\sigma_{i}^{\prime}-\frac{\sum_{j=1}^{n}\sigma_{j}^{\prime}}{m})^{2}}}$ (3)
We define ``scale'' as the ratio of the largest singular value and smallest singular value.
$\mbox{scale}=\frac{\sigma_{max}}{\sigma_{min}}$ (4)
Naturally, when the ``std'' is getting larger, the prompt matrix is getting closer to a low-rank form. The same holds true for the ``scale''.
Figure 2: Curves for the generalization ability (represented by ``new acc'' here) and the singular value distribution of prompt matrix. When ``new acc'' increases, ``std'' and ``scale'' also increase, i.e. the prompt matrix is getting closer to a low-rank form.
Table 1: Spearman correlation coefficients. Clearly, accuracy in the new classes is strongly correlated with std/scale of the singular values of the prompt matrix.
SUN397 DTD
std scale std scale
new acc 0.967 0.970 0.907 0.782
We show the result of each training epoch on base-to-new generalization setting in Fig. 2. The training curves are on two datasets, SUN397 [46] and DTD [4], respectively. Here we only show the test accuracy in the unseen new classes, which could measure the generalization ability of the model. For better visualization, we entirely scale up ``std'' to 8x and 6x on SUN397 and DTD respectively to match the value of ``scale'', so we could put them under the same axis in the figure. We also compute the Spearman correlation coefficient in Tab. 1, which measures the statistical dependence between the rankings of two variables. Seen from Figs. 2 and 1, the generalization ability of the vision-language model is highly correlated with ``std'' and ``scale'', e.g. the Spearman correlation coefficient of ``std'' and ``new acc'' is 0.967 on SUN397.
In summary, during training, the adaptation ability of the model represented by accuracy on the base classes is consistently improved as we have demonstrated in Fig. 5. The generalization ability of the model is improved in the first few epochs as well. However, as the training goes on, the generalization ability starts dropping, which indicates that the model suffers from overfitting and catastrophic forgetting problems. Seen from Fig. 2, the fluctuation trend of ``new acc'' is highly consistent with ``std'' and ``scale''. Naturally, we can infer that the low-rank form of the prompt matrix is beneficial to the generalization ability of the vision-language models. Motivated by such observation and inference, we propose Re-parameterized Low-rank Prompts (RLP) to get better generalization ability. Moreover, low-rank prompts naturally contain fewer parameters compared with classic ones, and thus we can also earn efficiency improvements.
@@ 3.1 Vision-Language Models
Recently, large-scale vision-language models have shown very competitive performance in various tasks. Classic works [37, 17, 51, 48, 49] learn the multi-modal representation by a self-supervised manner on a large amount of image-text pairs. The representative work CLIP [37] is a milestone, which aligns the vision representation and language representation by contrastive learning and shows excellent performance. A well-trained vision-language model is a great treasure, which could largely facilitate the development of various related fields. There have been successful applications of such strong models on few-shot recognition [55, 54], detection [39, 32, 9, 50] and segmentation [26, 38, 6, 31]. For video data, there are also works on video classification [36] and video understanding [20].
@@ 3.2 Prompt Tuning
Originating from natural language processing, prompts are first introduced as a fixed template [41], e.g. a photo of a _, which is hand-crafted and fixed. Later, a series of methods [27, 25, 29, 42, 28, 19] are proposed to make such prompts tunable and be optimized during adaptation. Prompt tuning could adaptively narrow the gap between pre-trained representations and downstream tasks, significantly facilitating the fine-tuning process. Representative prompt tuning methods would add tunable virtual tokens, i.e. prompts, along with the semantic tokens as inputs of the model. All of the tokens are processed together to get text embeddings first and then sent to the feature encoder. Witnessing the success of prompting language models, researchers design prompts [18, 52] for visual models in a similar way. In vision-language field, there are several explorations as well. Bahng et al. [1] adopts prompt tuning merely on the image encoder. CoOp [54] uses tunable text prompts to replace the fixed template in CLIP [37]. CoCoOp [55] utilizes image feature to instruct the optimization of the tunable text prompts in CoOp.  [21, 24] simultaneously optimize image and text prompts and establish extra connections between different modals.  [22, 47, 3, 53] integrate strong regularization modules or losses into prompt tuning to diminish the overfitting and catastrophic forgetting problem. For better downstream accuracy, researchers design more and more complicated methods, accompanied by inefficiency. To solve the problem, we propose Re-parameterized Low-rank Prompt (RLP) to take the place of classic prompts, which can largely decrease the number of tunable parameters and further enhance the model's generalization ability. Notice that though becoming more complex, existing methods are still refined on a common fundamental basis, i.e. prompt tuning. Such a common basis guarantees that RLP could be easily and smoothly integrated into most of the off-the-shelf methods besides individually applied.
@@ 4 Methodology
4.1 A Review of Prompt Tuning for CLIP
In this subsection, we first make a brief review of CLIP [37] which is used as our foundational model. CLIP consists of a text encoder $\mathcal{L}$ and an image encoder $\mathcal{V}$. Typically, $\mathcal{L}$ is a language transformer, while $\mathcal{V}$ can be a convolutional neural network or a vision transformer. In this paper, we follows  [54, 55] to use a ViT-B/16 [7] as the image encoder $\mathcal{V}$ unless specifically mentioned. We will make a review of how to prompt a CLIP for downstream image recognition in the following paragraphs.
Text Encoder Suppose there are $M$ layers in the text encoder. For $k$-th layer $\mathcal{L}_{k}$, the inputs are a series of prompt tokens $P^{l}_{k-1}$ and a [CLS] token $c^{l}_{k-1}$, and the outputs are $P^{l}_{k}$ and $c^{l}_{k}$. The inputs of the first layer $P^{l}_{0}$ and $c^{l}_{0}$ are exactly the word embeddings of the prompts along with the label, e.g. ``A photo of a [CLS]'' or just some randomly initialized vectors. Formally, we have $P^{l}_{k}\in\mathbb{R}^{n^{l}\times d^{l}}$ and $c^{l}_{k}\in\mathbb{R}^{d^{l}}$, where $n^{l}$ denotes the text prompts' length and $d^{l}$ denotes the dimension of word embedding. $\forall 1\leq k\leq M$, we have
$[P^{l}_{k},c^{l}_{k}]=\mathcal{L}_{k}([P^{l}_{k-1},c^{l}_{k-1}])$ (5)
Image Encoder Suppose there are $N$ layers in the image encoder. For $k$-th layer $\mathcal{V}_{k}$, the inputs are a series of image patch tokens $I_{k-1}$, a classification token $c^{v}_{k-1}$ and prompt tokens $P^{v}_{k-1}$, and the outputs are $I_{k}$, $c^{v}_{k}$ and $P^{v}_{k}$. The inputs of the first layer $I_{0}$ and $c^{v}_{0}$ are exactly the patch embeddings of the input image and the pre-trained class token. $P^{v}_{0}$ is randomly initialized in general. Formally, we have $I_{k}\in\mathbb{R}^{p\times d^{v}}$, $c^{v}_{k}\in\mathbb{R}^{d^{v}}$ and $P^{v}_{k}\in\mathbb{R}^{n^{v}\times d^{v}}$, where $p$ denotes the number of image patches and $d^{v}$ denotes the dimension of visual embedding. $\forall 1\leq k\leq N$, we have
$[P^{v}_{k},c^{v}_{k},I_{k}]=\mathcal{V}_{k}([P^{v}_{k-1},c^{v}_{k-1},I_{k-1}])$ (6)
The output feature of the image encoder is $f^{v}=c^{v}_{N}$.
Prediction CLIP can be easily used for image classification. Suppose there are $C$ classes, and $\{f^{l}_{c}\}_{c=1}^{C}$ are the corresponding text features. The probability of label $y$ is
$p(y|f^{v})=\frac{\mbox{exp}(\mbox{sim}(f^{v},f^{l}_{y})/\tau)}{\sum_{c=1}^{C}\mbox{exp}(\mbox{sim}(f^{v},f^{l}_{c})/\tau)}$ (7)
where sim(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denotes cosine similarity function and $\tau$ is temperature. The final prediction is $\hat{z}=\mathop{\arg\max}\limits_{1\leq y\leq C}(p(y|f^{v}))$.
It is worth noting that some researchers adopt a deeper manner [18, 21] to organize the prompts. They directly add and tune the prompt in each layer in the feature encoder, instead of inheriting the output prompt calculated by the last encoder, i.e. a forward pass becomes
[_$[\_,c^{l}_{k}]=\mathcal{L}_{k}([P^{l}_{k-1},c^{l}_{k-1}])$ (8)
and
[_$[\_,c^{v}_{k},I_{k}]=\mathcal{V}_{k}([P^{v}_{k-1},c^{v}_{k-1},I_{k-1}])$ (9)
Each $P^{l}$/$P^{v}$ contains tunable parameters.
4.2 Re-parameterized Low-rank Prompts
4.2.1 Low-rank Decomposition
Motivated by the observation in Sec. 2, we re-parameterize the origin full-rank prompt matrix into a low-rank form in adaptation for better generalizability and efficiency.
Formally, given a typical text or image prompt matrix $P\in\mathbb{R}^{n\times d}$, we take inspiration from LoRA [16] to create two low-dimentional matrices and use their product as the final low-rank prompt. Suppose the intrinsic rank is $r$, then we first randomly initialize $P_{A}\in\mathbb{R}^{n\times r}$ and $P_{B}\in\mathbb{R}^{r\times d}$. The low-rank prompt $P_{lr}$ is
$P_{lr}=P_{A}P_{B}$ (10)
Here $P_{lr}$ is in the same shape with normal full rank prompt $P$. As a result, we could do a low-rank prompt tuning by sending $P_{lr}$ rather than the original $P$ into the corresponding feature encoder as shown in Eqs. 5 and 6, and updating the parameters in $P_{A}$ and $P_{B}$ through the back-propagated gradients.
4.2.2 Full-rank Initialization
Figure 3: To take full advantage of some special initialization form of prompts, e.g. a hand-crafted template a photo of a for the text prompts, we introduce a concurrent full-rank prompt branch along with the proposed low-rank prompts. By turning off the gradient of the newly added branch, we start training from a promising initial point, and the total number of tunable parameters or stored parameters will not increase as well. The Dropout layer could effectively regularize the update of low-rank prompts and alleviate overfitting and catastrophic forgetting. Better still, Dropout is a lightweight non-parametric layer and turns out to be an Identity layer in inference, resulting in negligible cost.

### v3

# Towards Efficient Vision-Language Tuning: More Information Density, More Generalizability

@@ Abstract
With the advancement of large pre-trained vision-language models, effectively transferring the knowledge embedded within these foundational models to downstream tasks has become a pivotal topic, particularly in data-scarce environments. Recently, parameter-efficient fine-tuning approaches, especially prompt tuning, have garnered considerable attention. To better understand the nature of prompt tuning, we propose the concept of “Information Density” (ID) to indicate whether a matrix strongly belongs to certain feature spaces rather than being evenly distributed across various feature spaces. We suppose a higher ID with strong bias across some feature spaces naturally leads to excellent robustness and stability. Our research, inspired by the observation that generalizability is closely linked to the information density of the prompt matrix, introduces the Dense Information Prompt (DIP). DIP aims to enhance information density to improve generalization. Furthermore, DIP significantly reduces the number of tunable parameters and the requisite storage space, making it particularly advantageous in resource-constrained settings. Comprehensive experiments substantiate the superiority of DIP. Notably, DIP surpasses the latest state-of-the-art methods by a substantial margin with an exceptionally small parameter count. Across a range of tasks spanning 11 datasets, DIP improves the average downstream accuracy of classic prompt tuning by up to 5.76% using merely 0.5K parameters.
33footnotetext: Equal contribution44footnotetext: Corresponding authors
@@ 1 Introduction
However, traditional fine-tuning is not a good choice to adapt vision-language models. Simply fine-tuning all the parameters can easily cause the model to overfit because the huge number of parameters bring redundant non-essential information. The huge training and storage cost is also an intractable problem. In the context of our study, we introduce the concept of “information density”. Much like the rank of a matrix in linear algebra, which represents the maximum number of linearly independent rows or columns in the matrix, “information density” represents the maximum amount of essential and non-redundant information that the model can extract from the downstream task. Just as a matrix with a higher rank possesses more unique information, a model with high information density can acquire more essential and general information from the downstream task by fewer parameters, even with a smaller dataset. Our goal is increasing the information density and thus using the fewest but most essential parameters to finish generalization, without causing catastrophic forgetting or overfitting to the small dataset.
As the concept of information density can be functionally analogous to the rank of a matrix, we decided to use properties related to rank to quantify information density. Specifically, we take a full-rank matrix and decompose it using Singular Value Decomposition (SVD) to obtain a matrix of singular values, using the properties of these singular values to define and quantize information density. We found that this definition of information density is highly correlated with the model’s generalization performance (Spearman correlation coefficient $>$ 0.9). Therefore, we propose DIP, aiming to enhance the model’s generalization ability by increasing information density. Additionally, due to the increased information carried by each parameter unit, our approach can significantly reduce the required number of parameters.
In Section 3, we will show our finding about the strong correlation between generalization capability and information density of the prompt matrix in Fig. 1. Inspired by such observation, we propose Dense Information Prompt (DIP) for effective and efficient adaptation, well adapting models under an extremely small number of parameters where nobody has explored before. There are several good advantages of DIP:
  • Efficiency and Effectiveness To our knowledge, we are the first one to reach comparable or even better performance with state-of-the-art methods using such an extremely small number of parameters, i.e. 516.
  • Simplicity Replacing the classic prompt by DIP just needs to modify several lines of code. No special loss nor extra module is introduced.
  • Robustness DIP is relatively capable of anti-disturbance. As in Tab. 3, DIP could maximally reserve its knowledge from domain shift.
  • Plug and Play For any classic model, DIP only replaces the prompts, enabling us to plug DIP into most of the existing methods fruitfully.
@@
  • We propose a new concept “information density”, give its definition and well-quantize the concept. We further find the strong correlation between generalizability and information density in Fig. 1, and thus propose to use DIP to increase information density for better generalization capability.
  • We propose a novel initialization method and integrate a lightweight regularization module to further improve the performance of Dense Information Prompt tuning without introducing any extra parameters and inference cost.
@@
  • We conduct extensive experiments and show the fantastic effectiveness and efficiency of DIP. In base-to-new generalization, domain generalization, cross-dataset transfer and few-shot learning settings, DIP consistently reaches very competitive results, surpassing many state-of-the-art tuning methods though we use fewer parameters.
@@ 2.1 Vision-Language Models
Recently, large-scale vision-language models have shown very competitive performance in various tasks. Classic works [1, 2, 3, 4, 5] learn the multi-modal representation by a self-supervised manner on a large amount of image-text pairs. The representative work CLIP [1] is a milestone, which aligns the vision representation and language representation by contrastive learning and shows excellent performance. A well-trained vision-language model is a great treasure, which could largely facilitate the development of many fields. There have been successful applications of such strong models on few-shot recognition [6, 7], detection [8, 9, 10, 11] and segmentation [12, 13, 14, 15]. For video data, there are also works on video classification [16] and video understanding [17].
@@ 2.2 Prompt Tuning
Prompt tuning is one of the most popular methods to tune models in downstream tasks with excellent efficiency. Originating from natural language processing, prompts are first introduced as a fixed template [18], e.g. a photo of a _, which is hand-crafted and fixed. Later, a series of methods [19, 20, 21, 22, 23, 24] are proposed to make such prompts tunable and be optimized during adaptation. Prompt tuning could adaptively narrow the gap between pre-trained representations and downstream tasks, significantly facilitating the fine-tuning process. Representative prompt tuning methods would add tunable virtual tokens, i.e. prompts, along with the semantic tokens as inputs of the model. All of the tokens are processed together to get text embeddings first and then sent to the feature encoder. Witnessing the success of prompting language models, researchers design prompts [25, 26] for visual models in a similar way. In vision-language field, there are several explorations as well. Bahng et al. [27] adopts prompt tuning merely on the image encoder. CoOp [7] uses tunable text prompts to replace the fixed template in CLIP [1]. CoCoOp [6] utilizes image feature to instruct the optimization of the tunable text prompts in CoOp.  [28, 29] simultaneously optimize image and text prompts and establish extra connections between different modals.  [30, 31, 32, 33, 34] integrate strong regularization modules or losses into prompt tuning to diminish the overfitting and catastrophic forgetting problem. For better downstream accuracy, researchers design more and more complicated methods, accompanied by inefficiency. To solve the problem, we propose Dense Information Prompt (DIP) to take the place of classic prompts, which can largely decrease the number of tunable parameters and further enhance the model’s generalization ability. Notice that though becoming more complex, existing methods are still refined on a common fundamental basis, i.e. prompt tuning. Such a common basis guarantees that DIP could be easily and smoothly integrated into most of the off-the-shelf methods besides individually applied. Besides prompt-based methods, there are also many other works to acquire storage efficiency [35, 36, 37, 38, 26, 39], inference efficiency [40, 41, 42, 43, 44, 45, 46, 47, 48] and data efficiency [49, 50] during downstream tuning.
3 Relationship between Information Density and Generalizability
In this section, we will start from reviewing a classic prompt tuning pipeline CoOp [7] on CLIP in Section 3.1, and propose a new concept “Information Density” and analyze its relationship with generalizability in Section 3.2.
3.1 A Review of Prompt Tuning for CLIP
CLIP consists of a text encoder $\mathcal{L}$ and an image encoder $\mathcal{V}$. Typically, $\mathcal{L}$ is a language transformer, while $\mathcal{V}$ can be a convolutional neural network or a vision transformer. In this paper, we follows  [7, 6] to use a ViT-B/16 [51] as the image encoder $\mathcal{V}$ unless specifically mentioned. We start by making a review of how to prompt a CLIP for prediction in the following paragraphs.
Text Encoder Suppose there are $M$ layers in the text encoder. For $k$-th layer $\mathcal{L}_{k}$, the inputs are a series of prompt tokens $P^{l}_{k-1}$ and a [CLS] token $c^{l}_{k-1}$, and the outputs are $P^{l}_{k}$ and $c^{l}_{k}$. The inputs of the first layer $P^{l}_{0}$ and $c^{l}_{0}$ are exactly the word embeddings of the prompts along with the label, e.g.A photo of a [CLS]” or just some randomly initialized vectors. Formally, we have $P^{l}_{k}\in\mathbb{R}^{n^{l}\times d^{l}}$ and $c^{l}_{k}\in\mathbb{R}^{d^{l}}$, where $n^{l}$ denotes the text prompts’ length and $d^{l}$ denotes the dimension of word embedding. $\forall 1\leq k\leq M$, we have $[P^{l}_{k},c^{l}_{k}]=\mathcal{L}_{k}([P^{l}_{k-1},c^{l}_{k-1}])$.
Image Encoder Suppose there are $N$ layers in the image encoder. For $k$-th layer $\mathcal{V}_{k}$, the inputs are image patch tokens $I_{k-1}$, a classification token $c^{v}_{k-1}$ and prompt tokens $P^{v}_{k-1}$, and the outputs are $I_{k}$, $c^{v}_{k}$ and $P^{v}_{k}$. The inputs of the first layer $I_{0}$ and $c^{v}_{0}$ are exactly the patch embeddings of the image and pre-trained class token. $P^{v}_{0}$ is randomly initialized in general. Formally, we have $I_{k}\in\mathbb{R}^{p\times d^{v}}$, $c^{v}_{k}\in\mathbb{R}^{d^{v}}$ and $P^{v}_{k}\in\mathbb{R}^{n^{v}\times d^{v}}$, where $p$ denotes the number of image patches and $d^{v}$ denotes the dimension of visual embedding. $\forall 1\leq k\leq N$, $[P^{v}_{k},c^{v}_{k},I_{k}]=\mathcal{V}_{k}([P^{v}_{k-1},c^{v}_{k-1},I_{k-1}])$. The output feature of the image encoder is $f^{v}=c^{v}_{N}$.
Prediction CLIP can be used for image classification. Suppose there are $C$ classes, and $\{f^{l}_{c}\}_{c=1}^{C}$ are the corresponding text features. Label $y$’s probability is $p(y|f^{v})=\frac{\mbox{exp}(\mbox{sim}(f^{v},f^{l}_{y})/\tau)}{\sum_{c=1}^{C}\mbox{exp}(\mbox{sim}(f^{v},f^{l}_{c})/\tau)}$ where sim(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denotes cosine similarity function and $\tau$ is temperature. The final prediction is $\hat{z}=\mathop{\arg\max}\limits_{1\leq y\leq C}(p(y|f^{v}))$.
It is worth noting that some researchers adopt a deeper manner [25, 28] to organize the prompts. They directly add and tune the prompt in each layer in the feature encoder, instead of inheriting the output prompt calculated by the last encoder, i.e. a forward pass becomes [_$[\_,c^{l}_{k}]=\mathcal{L}_{k}([P^{l}_{k-1},c^{l}_{k-1}])$ and [_$[\_,c^{v}_{k},I_{k}]=\mathcal{V}_{k}([P^{v}_{k-1},c^{v}_{k-1},I_{k-1}])$. Each $P^{l}$/$P^{v}$ contains tunable parameters.
Figure 1: Relationship between generalizability represented by the test accuracy on unseen classes during training and Information Density (ID). When generalizability increases, ID also increases. The Spearman correlation coefficient $\rho$ between generalizability and ID1/ID2 is very high, i.e. $\geq$ 0.9.
3.2 Information Density in Prompt Tuning
Here, we first provide precise definitions for “information density” to clearly convey our motivations. For typical parameter matrix like prompts $P\in\mathbb{R}^{n\times d}$(assume $n<d$), we can always rewrite such matrix into a combination of several orthogonal bases with different weights by singular value decomposition (SVD). Formally, $P=U\Sigma V^{T}=\sum_{i=1}^{d}\sigma_{i}u_{i}v_{i}^{T}$. Each $u_{i}v_{i}^{T}$ can span a unique feature space, and $P$ is a linear combination of these features. Typically, $\{\sigma_{i}\}_{i=1}^{n}$ are arranged in descending order.
[…]
Returning to our initial discussion, a core contribution of this paper is the hypothesis and verification that the parameter matrices, like prompts, follow the same ID-related laws during fine-tuning in downstream CLIP models. As is well-known, in transfer learning, the transfer of knowledge in a model depends on the updating of parameters. For CLIP models, the optimal solution is prompt tuning [7, 6, 28, 52]. We first hypothesize that the information density of the prompt matrix also represents its robustness and the strength of its intrinsic features. Therefore, greater information density should theoretically result in better generalizability throughout the prompt tuning process.
To verify this hypothesis, we conducted an experiment using a classic method, CoOp [7]. During training, we masked half of the classes, using data from only the other half for training, and performed singular value decomposition on the prompt matrices during the process. As shown in Fig. 1, for better visualization, ID1 is scaled up to 2x to be put under the same right axis with ID2. Unseen classes accuracy is improved in the first few iterations, but it starts dropping later, indicating overfitting and catastrophic forgetting. Importantly, the fluctuation trend of unseen classes accuracy is highly consistent with ID. We compute the Spearman correlation coefficient between unseen classes accuracy and ID1/ID2 in Fig. 1. Clearly, the first-order and second-order information densities of CoOp on the SUN397 [53] and DTD [54] datasets exhibit a very strong correlation with the accuracy on unseen classes (i.e., generalizability), with Spearman correlation coefficients greater than 0.9. This demonstrates that our hypothesis is correct.
In the following Section 4, we will show how we can leverage such correlation between generalizability and ID to boost CLIP’s downstream performance.
@@ 4 Methodology
4.1 Dense Information Prompt
4.1.1 Algorithms for increasing information density
In this part, we aim to find an effective algorithm to increase information density to enhance CLIP’s downstream generalization performance. Motivated by the observation in Section 3, we propose three optional algorithms.
[…]
We do a quick verification on several datasets used in the base-to-new generalization setting and show the results in Tab. 1. Clearly, from Alg. 1 to Alg. 3, the intensity of constraint is becoming stronger, resulting in more accuracy loss on base classes while enjoying more accuracy gain on new classes. In addition to the good performance, Alg. 3 has another advantages: it’s quite simple and easy to implement with faster training speed. Therefore, we choose Alg. 3 to implement our DIP. However, unlike LoRA who merely uses a similar low-rank product as a supplementary addition for the output of main branch, Alg. 3 directly outputs the product for usage, leading to some unique difficulties and challenges. We will show the problems behind such approximation and our solutions in the next sections.
Table 1: A quick check for different implementations of DIP.
base new H
w/o any 82.69 63.22 71.66
Alg. 1 81.21 67.58 73.77
Alg. 2 80.39 69.85 74.75
Alg. 3 79.91 71.48 75.46
4.1.2 Special Initialization